This week John Good, a mechanical engineer in Austin, is back out to run the 97 point test to characterize the shape of the sphere that the tracker traces out above the primary mirror. Several weeks ago we attempted to run this test but found that the number of measurements we were taking per point was to many and we were only able to get a 21 point test. This time the tracker is running more reliably and we are going to take fewer measurements per point. This test will be conducted at night and we expect it will take 3 nights.
We no longer need the four mirrors that we removed from the array for laser tracker work. The optics team will begin to replace these last 4 mirrors, at which time we should have a fully populated array.
The mechanical contractors on site continue to work on the pipes in the vault. This week they will be putting insulation around the pipes. The dirt works contractors are working on the base of the piers which will hold the acid vapor from the strip and wash room and heat ducts from the Krabenhut.
In the past week and this week the contractors on site continue to work on their projects. Veliz Construction is half way done with the application of the silver tape to lower the emissivity of the control and services building next to the dome. We expect them to take about 2 weeks to finish the rest of the tape work if the afternoon rains keep their current pattern and stop their work. ARC Mechanical continue to populate the vault with freon, glycol and other lines as well as connecting up to the inside of the building. This work progresses slowly.
We have gotten the approval for the opto-mechanical technician position. If you have interest or know someone interested in working in the beautiful Davis Mountains and working with the optics of one of the largest telescopes in the world please have them apply to: job number 140715014211
Last week we had the metrology team out from Austin and reinstalled the VAT, tip-tilt camera and VAT on the telescope. This trip we had much greater success. We were able to get the system aligned and got a return from the DMI. We also were able to turn the VAT around and point at the sky. We got our first trajectory and followed the star we actually desired for 20 minutes in which time the star did not drift appreciably. We did have some problems getting trajectories to start so that was the focus of the work this week. After a few days of digging we found a variable that was defined in the wrong part of the code and with the variable moved to the right location we were able to run 8 trajectories one after another and also ran trajectories that simulated stars moving in the East and North (we had been working in the South were there is little rotation and the trajectories move more quickly). This is a big break through!
In the past week the remote thermal area contractors finished the work on the remote concrete pads and the trench between these pads and the vault. This week the contractors will work on the Al tape coating for the roof of the control and services building to reduce the emissivity and avoid the super cooling of the roof. The mechanical contractor will work on the conduit that lead into the building from the vault.
We have had the departure from the opto-mechanical team. We will start the hunt for a new technician as soon as HR approves our new job posting.
This week we will be installing the DMI, tip-tilt camera and VAT on the telescope. In a previous week the VAT and mounts were test fit so we hope that all of the components go on easily. The software team will also be coming out and we expect to turn the VAT towards the sky and have a first track and guide using the TCS software.
In the past week we have had a number of very strong thunder storms one of which blew out one of our 480 V contactors in the new tracker electronics. We found we don’t have a spare for this relatively inexpensive components so we are ordering a replacement and several spares. We hope to have this in place in the next day or so. Once back up we will continue to work on the TCS trajectory generation.
Next week we have some of the engineers and most of the software team coming out from Austin for a first attempt at running trajectories with metrology. The DMI is back from France and ready to be installed on the test mass along with the tip-tilt camera and the Video Alignment Telescope (VAT). This week we are just making sure all of the communications and power connections are in place for this set of tests.
We have four different contractors on site this week working on the Remote Thermal Area project. One is building the forms for the remote concrete pads, another is digging a trench between those pads and the vault which was completed last week, another is putting in ducting in the vault and the last is installing ducting inside the spectrograph room, lower and upper electrical rooms. The afternoon storms are slowing the work slightly.